OpenStudy (anonymous):

Calculus The following formula accurately models the relationship between the size of a certain type of a tumor and the amount of time that it has been growing: V(t)=400(1-e^ -.0024t)^3 where t is in months and V(t) is measured in cubic centimeters. Calculate the rate of change of tumor volume at 100 months.

OpenStudy (freckles):

have you differentiated yet?

OpenStudy (freckles):

@mthompson440 are you there?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

well, I have tried to differentiate but I must not be doing it right. V'(100)= 3* (-.0024)* (1-e^ -.24)^2 but that is not correct

OpenStudy (freckles):

\[V=400(1-e^{-0.0024t})^3\] this right for V?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

I'm going to start on the most outside I do notice your 400 constant multiple is missing right off the bat and you have changed -.0024 to -.24 but let's go through it slowly

OpenStudy (freckles):

oh you plug in a 100 :p

OpenStudy (anonymous):

yes )

OpenStudy (freckles):

\[V'=400 \cdot 3(1-e^{-0.0024 t})^2 \cdot (1-e^{-0.0024t})'\] so here I brought down constant multiply of 400 and applied chain rule to the ( )^3 part

OpenStudy (freckles):

now we nee to also different (1-exp(-0.0024t)

OpenStudy (freckles):

\[\frac{d}{dt}(1-e^{-0.0024 t})=\frac{d}{dt}(1)-\frac{d}{dt}e^{-0.0024t}=0-\frac{d}{dt}e^{-0.0024t } \\ -\frac{d}{dt}e^{-0.0024t}=?\] for this part you need chain rule also

OpenStudy (anonymous):

I thought that the 400 being a constant = 0 ?

OpenStudy (freckles):

400 is a constant but it is a constant multiple it does not stand alone

OpenStudy (freckles):

\[\frac{d}{dx}400=0 \\ \text{ but } \frac{d}{dx}400x=400\]

OpenStudy (anonymous):

= e^.0024t or e^.0024(100) or e^.24

OpenStudy (freckles):

can you differentiated what I asked you to differentiate above?

OpenStudy (freckles):

\[\frac{d}{dx}e^{ax} =(ax)'e^{ax}=ae^{ax}\]

OpenStudy (freckles):

so \[-\frac{d}{dt} e^{-0.0024 t}=?\]

OpenStudy (anonymous):

the two negative become positive - (-e^0.0024t)

OpenStudy (freckles):

yes but you are still forgetting something

OpenStudy (freckles):

\[-\frac{d}{dt} e^{-0.0024t}=-(-0.0024t)'e^{-0.0024t}=-(-0.0024)e^{-0.0024t}=0.0024e^{-0.0024t}\]

OpenStudy (anonymous):

- .0024e^-.0024t

OpenStudy (freckles):

\[V'=400 \cdot 3(1-e^{-0.0024 t})^2 \cdot (1-e^{-0.0024t})' \\ V'=1200(1-e^{-0.0024t})^2 \cdot 0.0024 e^{-0.0024 t}\]

OpenStudy (freckles):

now we can replace t with 100

OpenStudy (freckles):

we can do one more thing before do that

OpenStudy (freckles):

\[V'=2.88(1-e^{-0.0024t})^2 e^{-0.0024 t}\]

OpenStudy (freckles):

any questions?

OpenStudy (anonymous):

Thank you so much. I will have to sit here and digest this. But it all looks good. I would not have figured this out. Thank you so much

OpenStudy (freckles):

do you understand about the constant multiple part?

OpenStudy (freckles):

\[\frac{d}{dx}(c \cdot f(x)) = c \cdot \frac{d}{dx} f(x) \text{ this constant multiple rule }\]

OpenStudy (freckles):

you could also use product rule to come up with that

OpenStudy (freckles):

\[\frac{d}{dx} (c \cdot f(x)) \\ =f (x) \cdot \frac{d}{dx} c+c \cdot \frac{d}{dx} f(x) \\ =f(x) \cdot 0+ c \cdot \frac{d}{dx}f(x) \\ =0+c \cdot \frac{d}{dx}f(x) \\ =c \cdot \frac{d}{dx}f(x)\]

OpenStudy (freckles):

400 being the c here

OpenStudy (freckles):

of course we could use the rule again in this same function because we have a bit of chain rule going on